OpenStudy (anonymous):

Simplify 0.3 + 1 3/4

OpenStudy (anonymous):

0.3 + 1 + 0.75

OpenStudy (anonymous):

2.05

OpenStudy (anonymous):

into a fraction

OpenStudy (anonymous):

it needs to be made into a fraction guys

OpenStudy (anonymous):

2 + 1/20= 41/20

OpenStudy (diyadiya):

1 3/4 canbe canged into \[1 \frac{3}{4}= \frac{1 \times 4 + 3 }{4}= \frac{7}{4}\] 0.3 = 3/10 Now can you add \[\frac{3}{10} + \frac{7}{4}\]

OpenStudy (anonymous):

i got 41 over 20........but thats not the answer

OpenStudy (anonymous):

@Diyadiya

OpenStudy (anonymous):

then it shud be 82/40

OpenStudy (diyadiya):

You're Right! :) @daniellemmcqueen

OpenStudy (diyadiya):

what are your choices?

OpenStudy (anonymous):

A. 2 1/2 B. 1 3/7 C. 4 3/4 D. 2 1/20 @Diyadiya

OpenStudy (diyadiya):

Ok so you have to convert this to mixed fraction

OpenStudy (anonymous):

Mhm!

OpenStudy (anonymous):

Help? @Diyadiya

OpenStudy (diyadiya):

Sorrrrry i made a mistake